Hi I need the answer please :) I am using half of my questons in this one box I will submit my other three in another question and thats it becasue it will b unfair 2 others
Find the hypotenuse of a right triangle with legs measuring 5.2 cm and 7.8 cm respectively. Round your answer to the nearest tenth of a centimeter.
3.6 cm
5.8 cm
18.5 cm
9.4 cm
16.5 cm
I also need this answered :)
Sam hiked 48 miles directly east and then 36 miles directly south. Find the shortest distance in miles from the point where he ended up to his starting point.
72 miles
60 miles
48 miles
36 miles
24 miles
This one 2
The screen on Jeff's computer monitor has a length of 15.2 and a width of 11.5 in. Find the measure of the diagonal for Jeff's monitor.
19 inches
24 inches
28 inches
17 inches
13 inches
: Hi I need the answer please :) I am using half of my questons in this one box I will submit my other three in another question and thats it becasue it will b unfair 2 others
Find the hypotenuse of a right triangle with legs measuring 5.2 cm and 7.8 cm respectively. Round your answer to the nearest tenth of a centimeter.
3.6 cm
5.8 cm
18.5 cm
9.4 cm
16.5 cm
I also need this answered :)
Sam hiked 48 miles directly east and then 36 miles directly south. Find the shortest distance in miles from the point where he ended up to his starting point.
72 miles
60 miles
48 miles
36 miles
24 miles
This one 2
The screen on Jeff's computer monitor has a length of 15.2 and a width of 11.5 in. Find the measure of the diagonal for Jeff's monitor.
19 inches
24 inches
28 inches
17 inches
13 inches

Answer by nerdybill(1040)

All three questions relates to the application of Pythagorean theorem -- given a "right triangle" (a triangle where ONE of the three internal angles measures 90 degrees) the side "opposite" the 90 deg internal angle is the "hypothenuse".
.
hypontenuse^2 = side_a^2 + side_b^2
.
Find the hypotenuse of a right triangle with legs measuring 5.2 cm and 7.8 cm respectively. Round your answer to the nearest tenth of a centimeter.
hypotenuse = sqrt(5.2^2 + 7.8^2)
hypotenuse = sqrt(27.04 + 60.84)
hypotenuse = sqrt(87.88)
hypotenuse = 9.4 cm
.
Sam hiked 48 miles directly east and then 36 miles directly south. Find the shortest distance in miles from the point where he ended up to his starting point.
hypotenuse = sqrt(48^2 + 36^2)
hypotenuse = sqrt(2304 + 1296)
hypotenuse = sqrt(3600)
hypotenuse = 60 miles
.
The screen on Jeff's computer monitor has a length of 15.2 and a width of 11.5 in. Find the measure of the diagonal for Jeff's monitor.
hypotenuse = sqrt(15.2^2 + 11.5^2)
hypotenuse = sqrt(231.04 + 132.25)
hypotenuse = sqrt(363.29)
hypotenuse = 19.1 inches